# Bastani, Bastani, and McLaughlin

### v2

@@
There has been a surge of recent interest in automatically learning policies to target treatment decisions based on rich individual covariates. A common approach is to train a machine learning model to predict counterfactual outcomes, and then select the policy that optimizes the predicted objective value. In addition, practitioners also want confidence that the learned policy has better performance than the incumbent policy according to downstream policy evaluation. However, due to the winner’s curse—an issue where the policy optimization procedure exploits prediction errors rather than finding actual improvements—predicted performance improvements are often not substantiated by downstream policy optimization. To address this challenge, we propose a novel strategy called inference-aware policy optimization, which modifies policy optimization to account for how the policy will be evaluated downstream. Specifically, it optimizes not only for the estimated objective value, but also for the chances that the policy will be statistically significantly better than the observational policy used to collect data. We mathematically characterize the Pareto frontier of policies according to the tradeoff of these two goals. Based on our characterization, we design a policy optimization algorithm that uses machine learning to predict counterfactual outcomes, and then plugs in these predictions to estimate the Pareto frontier; then, the decision-maker can select the policy that optimizes their desired tradeoff, after which policy evaluation can be performed on the test set as usual. Finally, we perform simulations to illustrate the effectiveness of our methodology.
@@ 1 Introduction
Much of the early work on policy evaluation assumes that a policy is fixed prior to evaluation (Rosenbaum and Rubin 1983, 1984, Bang and Robins 2005). Due to increasing interest in policy optimization, recent work has proposed methodologies for correcting for the optimistic bias arising from the winner’s curse (Gupta et al. 2024, Andrews et al. 2024, Zrnic and Fithian 2024, Shchetkina and Berman 2024, Zhang et al. 2025). However, they only correct for the evaluation error—if the learned policy is ineffective, they cannot help compute a more effective policy. Furthermore, practitioners often want evidence that the learned policy actually increases performance—e.g., by obtaining statistically significant performance improvement compared to the incumbent policy (which we call the observational policy). Thus, a critical remaining question is how to identify a policy that improves performance according to a rigorous downstream evaluation methodology. Addressing this challenge requires rethinking the entire policy evaluation and optimization pipeline—our key insight is that policy optimization should account for how the policy will be evaluated downstream, which we call inference-aware policy optimization.
We formalize the inference-aware policy optimization problem and mathematically characterize the space of solutions to this problem. We focus on a simple policy evaluation methodology using inverse propensity weighting (IPW), which evaluates a learned policy by reweighting test outcomes to appear as if they were collected under the learned policy (Rosenbaum and Rubin 1983); we restrict the policy optimization algorithm to only see training data and test covariates, which suffices to avoid the winner’s curse during evaluation (Shchetkina and Berman 2024). Our goal is to design a policy optimization algorithm that aims to achieve statistically significant performance improvement. Specifically, we aim to compute policies that trade off (1) the expected objective value according to the downstream IPW estimate, and (2) the expected $z$-score of the IPW estimate used to assess whether the learned policy is statistically significantly better than the observational policy. Intuitively, if the learned and observational policies are quite different, then the IPW estimator will downweight a large part of the test dataset when evaluating the proposed policy (Li et al. 2011), thereby effectively reducing sample size and failing to establish statistically significant performance improvement. Thus, to be conservative, we want to keep the learned policy closer to the random policy used to collect data, avoiding large propensity weights that can result in high variance.
@@
Using this characterization, we show that the policy with the largest expected $z$-score is close to the observational policy, making a small but confident improvement. Specifically, it differs from the observational policy for all subjects where an expected improvement can be made, with larger differences when the counterfactual means of treatments differ more and their related variances are small. The step size reaches the point where some treatment option is eliminated for some unit. Moving along the frontier from this initial solution, the tradeoff between expected objective and statistical significance becomes strict—i.e., any strict increase in expected objective requires a strict decrease in statistical significance. Furthermore, we show that treatment options are monotonically eliminated from consideration—i.e., once a treatment probability becomes zero, it stays at zero for the rest of the frontier.
Next, we show how our theoretical results can be used to derive an inference-aware policy optimization algorithm. The main barrier is that solution to our convex program is specialized to assigning treatments to our test units instead of being a generalizable policy that can be applied to any unit. However, we note that the policy assigns treatments based purely on the mean and standard deviation of the predicted counterfactual outcomes; thus, we can straightforwardly devise a generalizable policy from our solution. A full policy optimization and evaluation pipeline consists of the following steps: (1) train a predictive model on the training dataset, (2) compute the Pareto frontier using our characterization and select a policy along this frontier, and (3) evaluate the efficacy of this policy using IPW. We implement this pipeline and compare it to several baseline approaches in a simulation study, demonstrating that it can effectively compute the Pareto frontier.
@@ 1.1 Related Work
Our approach builds on the causal model introduced by Rubin (1974). Specifically, the key challenge with policy evaluation is the need to estimate counterfactual outcomes—i.e., for each treatment unit (e.g., an individual), we only observe the outcome for the treatment that is assigned to that unit by the observational policy (called the factual outcome), not the outcomes for other treatments (called counterfactual outcomes). Then, the policy evaluation problem is to estimate the average treatment effect (ATE) of one treatment compared to another (usually the control. This problem does not directly consider targeted interventions, where the treatment is targeted based on unit covariates instead of being constant across all units. However, when a specific candidate targeting policy is given, then the problem of policy evaluation can be reduced to this one by considering the policy to be the intervention, and evaluating the efficacy of that policy.
We focus on the setting where there are no unobserved confounders; for instance, this is the case in applications where data is collected from a known observational policy such as random assignment. In this setting, there are two strategies that can be applied for policy evaluation. The first (called the direct method in Dudík et al. (2011)) is to train a predictive model to estimate counterfactual outcomes, and use it to evaluate the candidate policy. This approach assumes that the predictive model is accurate, which is typically not true in practice. An alternative strategy is inverse propensity weighting (Rosenbaum and Rubin 1983), which uses importance weights to correct for the shift from the observational policy to the candidate policy. This approach requires that the inverse propensity weights are known (e.g., this is the case when the observational policy is random assignment); otherwise, they must be estimated. One approach to further improve on importance weighting is doubly robust estimation (Robins et al. 1994), which subtracts a prediction of the treatment outcome from the realized outcome to normalize it, resulting in the augmented IPW (AIPW) estimator. When the predicted outcomes are accurate, this strategy can significantly reduce variance compared to the IPW estimator. Doubly robust estimation is often used in settings where IPWs must be estimated, but it can reduce variance even when they are known. While we do not explicitly include the doubly robust correction in our main derivation, it can be straightforwardly incorporated into our framework.
@@
However, these strategies all suffer from the winner’s curse—even if policy evaluation is unbiased for a single candidate policy, the objective value of the learned policy is optimistically biased. These approaches avoid this issue by using sample splitting—i.e., optimize the objective on an optimization set, and then perform policy evaluation (e.g., using IPW) on a held-out test set (Shchetkina and Berman 2024) (the policy optimization procedure can use the test covariates; only treatment assignments and outcomes need to be held out). More sophisticated approaches have recently been proposed to avoid the reduction in sample complexity from sample splitting (Gupta et al. 2024, Andrews et al. 2024, Zrnic and Fithian 2024, Zhang et al. 2025). However, all of these approaches can only correct for optimistic bias in policy evaluation; they cannot select a policy that is more likely to achieve statistically significant policy improvement, and often fail to do so in practice.
There are two existing strategies that can be viewed as performing policy optimization to improve statistical significance. The first is treatment pooling, which tries to pool treatment options to reduce the size of the policy class in consideration (Banerjee et al. 2025, Ma et al. 2022, Spiess et al. 2023, Ladhania et al. 2023, Saito et al. 2024); a related approach is identifying promising subgroups to reduce the size of the policy space (Spiess et al. 2023). This procedure can be viewed as reducing variance in exchange for a small amount of bias; however, it does so in an ad hoc way and can only be applied to small policy classes.
@@ 1.2 Contributions
Our contributions are three-fold. First, we formalize the inference-aware policy optimization problem, where the goal is to obtain a policy with statistically significant policy improvement. Second, we perform a theoretical analysis on the tradeoff between expected performance and statistically significant policy improvement—specifically, we compute the entire Pareto frontier of policies trading off expected improvement and statistical significance of improvement. Finally, we propose a novel policy optimization algorithm based on this analysis; our policy optimization algorithm is the first one designed to enable practitioners to obtain statistically significant policy improvements. By selecting a desired policy along the Pareto frontier, practitioners can obtain the best policy at a desired level of confidence.
2 Characterizing the Pareto Frontier
We characterize the Pareto frontier of policies that jointly maximize the expected value and expected significance of an inverse propensity weighting (IPW) estimate of the policy improvement over the observational policy used to collect an observational dataset. Specifically, we generate a superset of the Pareto frontier using a convex relaxation whose solution can be described via a continuous piecewise linear parameterization of policies varying from the observational policy to the expectation maximizing policy. Then, we prune this superset to form the Pareto frontier.
@@ 2.1 Problem Formulation
Policy evaluation combines the Rubin causal model with inverse propensity weighting (IPW) to provide an unbiased estimate of an (unobserved) policies performance using data collected by an observational policy. Under the Rubin causal model, we cannot observe the counterfactual outcomes units would realize under treatments other than the one they were assigned by the observational policy, inhibiting our ability to estimate the performance of other policies. IPW forms an unbiased estimate of any policies performance by reweighting the outcomes of the observational policy so they appear as if generated by a different policy. As these estimates can be quite noisy, confidently believing any policy delivers an improvement over the observational policy requires passing a significance test. Thus, when proposing policies, practitioners must find high performing policies which they can show are significantly better than the observational policy.
Following the Rubin causal model (Rubin 1974), we consider a set of units $n\in[N]$ that each receive a treatment $t\in\{0,\ldots,K\}$, yielding a random real-valued outcome $Y_{n,t}$. We assume that each of these outcomes are generated independently upon treatment assignment without interference between units. We let $\mu_{n,t}=\mathbb{E}\left[Y_{n,t}\right]$ denote the mean and $\sigma^{2}_{n,t}=\mathrm{Var}(Y_{n,t})$ of denote the variance of these counterfactual outcomes. Each unit is assigned a treatment $T^{o}_{n}$ according to a known observational policy $\pi^{o}_{n}$ generating an observed outcome $Y_{n}^{o}=Y_{n,T_{n}^{o}}$. Each observational policy $\pi^{o}_{n}\in\Delta^{K}$ is an element of the standard $K$-simplex, and treatment assignments are independent random variables $T^{o}_{n}\sim\mathrm{Categorical}(\pi_{n}^{o})$. We assume each treatment is assigned to each unit with positive weight (i.e. positivity). We refer to the collection $\mathcal{O}=\{(\pi^{o}_{1},T^{o}_{1},Y^{o}_{1}),\ldots,(\pi^{o}_{N},T^{o}_{N},Y^{o}_{N})\}$ as the observational dataset and let $\pi^{o}_{n,t}$ be the probability unit $n$ is assigned treatment $t$.
We can use $\mathcal{O}$ to compare the performance of an alternate policy $\pi\in\left(\Delta^{K}\right)^{N}$ to $\pi^{o}$ using inverse propensity weighting (IPW) (Rosenbaum and Rubin 1983). Specifically, the difference in expected performance between $\pi$ and $\pi^{o}$ is given by the sample average treatment effect
@@
is the inverse propensity weight. As $\hat{\tau}(\pi)$ is only an estimate of the improvement of $\pi$ over $\pi^{o}$, a $z$-test needs to be applied to gain confidence that $\tau(\pi)>0$; we use a $z$-test since are assuming we know the variances (otherwise a $t$-test would be used), and we expect to use large sample sizes where the central limit theorem guarantees $\hat{\tau}(\pi)$ is normally distributed. The $z$-score for $\hat{\tau}(\pi)$ is
@@
is the standard error of $\hat{\tau}(\pi)$. Typically the standard error is unknown, so the sample approximation
$s(\pi)\approx\hat{s}(\pi)=\sqrt{\frac{1}{N(N-1)}\sum_{n=1}^{N}(Y_{n}^{o}(W_{n}-1)-\hat{\tau}(\pi))^{2}}$
based on $\mathcal{O}$ is used to estimate the $z$-score. The $z$-score is required to reach some level of significance $Z_{\min}$ before the policy is accepted.
Our goal is to construct a policy which improves upon the observational policy as much as possible while passing a significance test on the observational dataset. To do so we have access to the counterfactual outcome distributions of $Y_{n,t}$ and the observational policy $\pi^{o}$, but not the observed treatments or outcomes. We formalize this goal as a multi-criterion optimization problem between $\tau(\pi)$ and $\mathbb{E}[Z(\pi)]=\frac{\tau(\pi)}{s(\pi)}$ for which we seek the Pareto frontier. Note that for all $\pi$, $Z(\pi)$ should be normal with $\mathrm{Var}(Z(\pi))=1$, meaning that maximizing expected $z$-score should also maximize the probability of reaching $Z_{\min}$.
2.2 Approach
We seek to compute the Pareto frontier $\Pi_{\mathrm{Opt}}$ of policies that maximize some tradeoff between the expected improvement over the observational policy and the expected $z$-score when evaluated using IPW in conjunction with $\mathcal{O}$. First, we show that a superset of $\Pi_{\mathrm{Opt}}$ is given by a family of convex optimization problems $\Pi_{\Lambda}$. Then, we decompose the first order condition for this family of problems in order to characterize $\Pi_{\Lambda}$, which allows us to characterize $\Pi_{\mathrm{Opt}}$.
More precisely, our goal is to find the Pareto frontier of policies maximizing expected improvement $\tau(\pi)$ and expected $z$-score $\frac{\tau(\pi)}{s(\pi)}$—i.e., policies which cannot be strictly improved in both expected improvement and $z$-score:
$\Pi_{\text{Opt}}=\left\{\pi\in\left(\Delta^{K}\right)^{N}\biggm|\not\exists\ \pi^{\prime}\ \in\left(\Delta^{K}\right)^{N}\text{ s.t. }\tau(\pi^{\prime})\geq\tau(\pi)\wedge\frac{\tau(\pi^{\prime})}{s(\pi^{\prime})}\geq\frac{\tau(\pi)}{s(\pi)}\wedge(\tau(\pi^{\prime}),s(\pi^{\prime}))\neq(\tau(\pi),s(\pi))\right\}.$
This frontier is difficult to characterize directly as it involves solving a multi-criterion optimization problem in which one of the objectives is not convex. We start by characterizing a superset of $\Pi_{\text{Opt}}$. {repeatproposition} We have $\Pi_{\mathrm{Opt}}\subseteq\Pi_{\Lambda}$, where $\tau_{\text{max}}=\max_{\pi}\tau(\pi)$ and
$\Pi_{\Lambda}=\bigcup_{\lambda\in[0,\tau_{\text{max}}]}\underset{\left\{\pi\in\left(\Delta^{K}\right)^{N}\middle|\tau(\pi)=\lambda\right\}}{\mathrm{argmin}}s^{2}(\pi).$

### v3

@@
There has been a surge of recent interest in automatically learning policies to target treatment decisions based on rich individual covariates. In addition, practitioners want confidence that the learned policy has better performance than the incumbent policy according to downstream policy evaluation. However, due to the winner’s curse—an issue where the policy optimization procedure exploits prediction errors rather than finding actual improvements—predicted performance improvements are often not substantiated by downstream policy evaluation. To address this challenge, we propose a novel strategy called inference-aware policy optimization, which modifies policy optimization to account for how the policy will be evaluated downstream. Specifically, it optimizes not only for the estimated objective value, but also for the chances that the estimate of the policy’s improvement passes a significance test during downstream policy evaluation. We mathematically characterize the Pareto frontier of policies according to the tradeoff of these two goals. Based on our characterization, we design a policy optimization algorithm that estimates the Pareto frontier using machine learning models; then, the decision-maker can select the policy that optimizes their desired tradeoff, after which policy evaluation can be performed on the test set as usual. Finally, we perform simulations to illustrate the effectiveness of our methodology.  [^1] We gratefully acknowledge Katherine Milkman, Dena Gromet, and the Behavior Change for Good Initiative for providing access to the megastudy data used in Section 6.1 of our paper. Furthermore, we would like to thank Ron Berman, Dean Eckles, Vishal Gupta, Dean Knox, Daniel Rock, Neha Sharma, Hummy Song, Alp Sungu, Sophie Yu, and members of the Wharton AI & Analytics Initiative for their helpful discussion and feedback.
@@ 1 Introduction
Much of the early work on policy evaluation assumes that a policy is fixed prior to evaluation (Rosenbaum and Rubin 1983, 1984, Bang and Robins 2005). Due to increasing interest in policy optimization, recent work has proposed methodologies for correcting the optimistic bias arising from the winner’s curse (Gupta et al. 2024, Andrews et al. 2024, Zrnic and Fithian 2024, Xu et al. 2025). However, they only correct for the evaluation error—if the learned policy is ineffective, they cannot help compute a more effective policy. Furthermore, practitioners often want evidence that the learned policy actually increases performance—e.g., by obtaining a statistically significant estimate of the learned policy’s performance improvement over the incumbent policy (which we call the observational policy). Thus, a critical remaining question is how to identify a policy that improves performance according to a rigorous downstream evaluation methodology. Addressing this challenge requires rethinking the entire policy optimization pipeline—our key insight is that policy optimization should account for how the policy will be evaluated downstream, which we call inference-aware policy optimization.
We formalize the inference-aware policy optimization problem and mathematically characterize the space of solutions to this problem. We focus on a simple policy evaluation methodology using inverse propensity weighting (IPW), which evaluates a learned policy by reweighting test outcomes to appear as if they were collected under the learned policy (Rosenbaum and Rubin 1983); we restrict the policy optimization algorithm to only see training data and test covariates, which suffices to avoid the winner’s curse during evaluation. Our goal is to design a policy optimization algorithm that aims to achieve statistically significant performance improvement. Specifically, we aim to compute policies that trade off (1) the expected objective value according to the downstream IPW estimate, and (2) the expected $z$-score of the IPW estimate used to assess whether the learned policy is statistically significantly better than the observational policy. Intuitively, if the learned and observational policies are quite different, then the IPW estimator will downweight a large part of the test dataset when evaluating the proposed policy (Li et al. 2011), thereby effectively reducing sample size and failing to establish a statistically significant performance improvement. Thus, to be conservative, we want to keep the learned policy closer to the random policy used to collect data, avoiding large propensity weights that can result in high variance.
@@
Using this characterization, we show that the policy with the largest expected $z$-score is close to the observational policy, making a small but confident improvement. Specifically, it differs from the observational policy for all subjects where an expected improvement can be made, with larger differences when the counterfactual means of treatments differ more and their related variances are small. The step size reaches the point where some treatment option is eliminated for some unit. Moving along the frontier from this initial solution, the tradeoff between expected objective and statistical significance becomes strict—i.e., any strict increase in expected objective requires a strict decrease in statistical significance. Furthermore, we provide a simple closed form for the largest expected $z$-score, effectively allowing for power calculations that determine whether a test dataset has enough useful variation to evaluate any policy’s improvement as significant.
Next, we show how our theoretical results can be used to derive our algorithm for inference-aware policy optimization (IAPO). The main barrier is that the solution to our convex program is specialized to assigning treatments to our test units instead of being a generalizable policy that can be applied to any unit. However, we note that the policy assigns treatments based purely on the mean and standard deviation of the predicted counterfactual outcomes; thus, we can straightforwardly devise a generalizable policy from our solution. A full policy optimization and evaluation pipeline consists of the following steps: (1) train a predictive model on the training dataset, (2) compute the Pareto frontier using our characterization and select a policy along this frontier, and (3) evaluate the efficacy of this policy using IPW. We implement this pipeline in multiple simulation studies, demonstrating that it can effectively extract policies that deliver confident improvements in cases where naïve policy optimization approaches fail.
@@ 1.1 Related Work
Our approach builds on the causal model introduced by Rubin (1974). Specifically, the key challenge with policy evaluation is the need to estimate counterfactual outcomes—i.e., for each treatment unit (e.g., an individual), we only observe the outcome for the treatment that is assigned to that unit by the observational policy (called the factual outcome), not the outcomes for other treatments (called counterfactual outcomes). Then, the policy evaluation problem is to estimate the average treatment effect (ATE) of one treatment compared to another (usually the control). This problem does not directly consider targeted interventions, where the treatment is targeted based on unit covariates instead of being constant across all units. However, when a specific candidate targeting policy is given, then the problem of policy evaluation can be reduced to this one by considering the policy to be the intervention, and evaluating the efficacy of that policy.
We focus on the setting where there are no unobserved confounders; for instance, this is the case in applications where data is collected from a known observational policy such as random assignment. In this setting, there are two strategies that can be applied for policy evaluation. The first (called the direct method in Dudík et al. (2011)) is to train a predictive model to estimate counterfactual outcomes, and use it to evaluate the candidate policy. This approach assumes that the predictive model is accurate, which is typically not true in practice. An alternative strategy is inverse propensity weighting (Rosenbaum and Rubin 1983), which uses importance weights to correct for the shift from the observational policy to the candidate policy. This approach requires that the inverse propensity weights are known (e.g., this is the case when the observational policy is random assignment); otherwise, they must be estimated. One approach to further improve on importance weighting is doubly robust estimation (Robins et al. 1994), which subtracts a prediction of the treatment outcome from the realized outcome to normalize it, resulting in the augmented IPW (AIPW) estimator. When the predicted outcomes are accurate, this strategy can significantly reduce variance compared to the IPW estimator. Doubly robust estimation is often used in settings where weights must be estimated, but it can reduce variance even when they are known. While we do not explicitly include the doubly robust correction in our main derivation, it can be straightforwardly incorporated into our framework.
@@
However, these strategies all suffer from the winner’s curse—even if policy evaluation is unbiased for a single candidate policy, the objective value of the learned policy is optimistically biased. These approaches avoid this issue by using sample splitting—i.e., optimize the objective on an optimization set, and then perform policy evaluation (e.g., using IPW) on a held-out test set (the policy optimization procedure can use the test covariates; only treatment assignments and outcomes need to be held out). More sophisticated approaches have recently been proposed to avoid the reduction in sample complexity from sample splitting (Gupta et al. 2024, Andrews et al. 2024, Zrnic and Fithian 2024, Xu et al. 2025). However, all of these approaches can only correct for optimistic bias in policy evaluation; they cannot select a policy that is more likely to achieve statistically significant policy improvement, and often fail to do so.
There are two existing strategies that can be viewed as performing policy optimization to improve statistical significance. The first is treatment pooling, which tries to pool treatment options to reduce the size of the policy class in consideration (Banerjee et al. 2025, Ma et al. 2022, Ladhania et al. 2023, Saito et al. 2024); a related approach is identifying promising subgroups to reduce the size of the policy space (Spiess et al. 2023). This procedure can be viewed as reducing variance in exchange for a small amount of bias; however, it does so in an ad hoc way and can only be applied to small policy classes.
@@ 1.2 Contributions
Our contributions are three-fold. (1) We formalize the inference-aware policy optimization problem, where the goal is to obtain a policy with statistically significant policy improvement. (2) We perform a theoretical analysis on the tradeoff between expected performance and statistically significant policy improvement—specifically, we compute the entire Pareto frontier of policies trading off expected improvement and statistical significance of improvement. We find a number of interesting insights (summarized below). (3) We propose a novel policy optimization algorithm based on this analysis; our policy optimization algorithm is the first one designed to enable practitioners to obtain statistically significant policy improvements. By selecting a desired policy along the Pareto frontier, practitioners can obtain the best policy at a desired level of confidence.
Now, we summarize some of the key insights from our theoretical analysis:
  • Structure of the Pareto frontier: The Pareto frontier ranges between two policies, one maximizing statistical significance (specifically, the $z$-score) and one maximizing expected outcomes. Between the two, there is a strict tradeoff between $z$-score and expected outcomes.
  • Policies on the frontier: Notably, policies along the frontier are stochastic. Rather than assign units deterministically to treatments (as is the case with many existing methodologies), they assign treatment probabilities. These probabilities start from that of the observational policy, and move away until they bind to zero or one (this movement may be non-monotonic when there are multiple treatments). Intuitively, if a unit has high variance outcomes, then keeping its treatment probability close to that of the observational policy removes variance by effectively dropping it from the sample. This finding shows that policy learning algorithms must learn stochastic policies to improve statistical power.
  • Statistical power: The statistical power of the optimal $z$-score monotonically increases with the number of available units. However, the behavior of the $z$-score for the expectation-maximizing policy is very different. There can be a substantial gap between this $z$-score and the optimal one. Worse, it can scale to zero as the number of units increases, meaning increasing sample size can actually worsen significance. This finding requires somewhat pathological conditions, but even under reasonable conditions, it can scale non-monotonically with the number of units.
2 General Framework
Our goal is to characterize the Pareto frontier of policies that jointly maximize the expected value and expected significance of an inverse propensity weighting (IPW) estimate of the policy improvement over the observational policy used to collect an observational dataset. We begin by formalizing this problem (Section 2.1), and establishing the convex optimization framework we use to solve it (Section 2.2). In Section 3, we examine the special case of a single treatment (i.e., treatment vs. control group), enabling us to derive closed-form expressions for the Pareto frontier and obtain qualitative insights. Then, in Section 4, we analyze the general case of multiple treatments; while we can no longer obtain a closed-form representation of the frontier, we nevertheless show that our key insights about its structure continue to hold.
@@ 2.1 Problem Formulation
Policy evaluation combines the Rubin causal model with inverse propensity weighting (IPW) to provide an unbiased estimate of an (unobserved) policy’s performance using data collected by an observational policy. Under the Rubin causal model, we cannot observe the counterfactual outcomes units would realize under treatments other than the one they were assigned by the observational policy, inhibiting our ability to estimate the performance of other policies. IPW forms an unbiased estimate of any policies performance by reweighting the outcomes of the observational policy so they appear as if generated by a different policy. As these estimates can be quite noisy, confidently believing any policy delivers an improvement over the observational policy requires passing a significance test. Thus, when proposing policies, practitioners must find high performing policies which they can show are significantly better than the observational policy.
Following the Rubin causal model (Rubin 1974), we consider a set of units $n\in[N]$ that each receive a treatment $t\in\{0,\ldots,K\}$, yielding a random real-valued outcome $Y_{n,t}$. We assume that each of these outcomes are generated independently upon treatment assignment without interference between units. We let $\mu_{n,t}=\mathbb{E}\left[Y_{n,t}\right]$ denote the mean and $\sigma^{2}_{n,t}=\mathrm{Var}(Y_{n,t})$ denote the variance of these counterfactual outcomes. Each unit is assigned a treatment $T^{o}_{n}$ according to a known observational policy $\pi^{o}_{n}$ generating an observed outcome $Y_{n}^{o}=Y_{n,T_{n}^{o}}$. Each observational policy $\pi^{o}_{n}\in\Delta^{K}$ is an element of the standard $K$-simplex, and treatment assignments are independent random variables $T^{o}_{n}\sim\mathrm{Categorical}(\pi_{n}^{o})$. We assume each treatment is assigned to each unit with positive weight (i.e. positivity). We refer to the collection $\mathcal{O}=\{(\pi^{o}_{1},T^{o}_{1},Y^{o}_{1}),\ldots,(\pi^{o}_{N},T^{o}_{N},Y^{o}_{N})\}$ as the observational dataset and let $\pi^{o}_{n,t}$ be the probability unit $n$ is assigned treatment $t$.
We can use $\mathcal{O}$ to compare the performance of an alternate policy $\pi\in(\Delta^{K})^{N}$ to $\pi^{o}$ using inverse propensity weighting (IPW) (Rosenbaum and Rubin 1983). Specifically, the difference in expected performance between $\pi$ and $\pi^{o}$ is given by the sample average treatment effect
@@
is the inverse propensity weight. As $\hat{\tau}(\pi)$ is only an estimate of the improvement of $\pi$ over $\pi^{o}$, a $z$-test needs to be applied to gain confidence that $\tau(\pi)>0$; we use a $z$-test as we assume knowledge of the variances $\sigma^{2}$ (otherwise a $t$-test would be used), and we expect to use large sample sizes where the central limit theorem guarantees $\hat{\tau}(\pi)$ is normally distributed. The $z$-score for $\hat{\tau}(\pi)$ is
@@
is the standard error of $\hat{\tau}(\pi)$. [^2] Typically, the standard error is unknown, so we instead use a sample approximation $s(\pi)\approx\hat{s}(\pi)$ in conjunction with the $t$-score. For simplicity, we assume the standard error is known, so we can use the $z$-score. The policy is accepted if the $z$-score exceeds some level $Z_{\text{min}}$.
Our goal is to construct a policy which improves upon the observational policy as much as possible while passing a significance test on the observational dataset. To do so, we have access to the counterfactual outcome distributions of $Y_{n,t}$ and the observational policy $\pi^{o}$, but not the observed treatments or outcomes. Note that for all $\pi$, $Z(\pi)$ should be approximately normal with $\mathrm{Var}(Z(\pi))=1$, meaning that to maximize our probability of achieving statistical significance, it suffices to maximize the expected $z$-score. Thus, we formalize our goal as computing the Pareto frontier $\Pi_{\mathrm{Opt}}$ for a multi-objective optimization problem, where the objectives are (1) the expected policy improvement $\tau(\pi)$, and (2) the expected $z$-score $\mathbb{E}[Z(\pi)]=\frac{\tau(\pi)}{s(\pi)}$ for the IPW estimator used in conjunction with the observational dataset $\mathcal{O}$. Mathematically, our goal is to compute
$\Pi_{\text{Opt}}=\bigg\{\pi\in(\Delta^{K})^{N}\biggm|$ $\not\exists\ \pi^{\prime}\ \in(\Delta^{K})^{N}$
$\text{ s.t. }\tau(\pi^{\prime})\geq\tau(\pi)\wedge\frac{\tau(\pi^{\prime})}{s(\pi^{\prime})}\geq\frac{\tau(\pi)}{s(\pi)}\wedge(\tau(\pi^{\prime}),s(\pi^{\prime}))\neq(\tau(\pi),s(\pi))\bigg\},$
i.e., the Pareto frontier $\Pi_{\text{Opt}}$ consists of exactly the policies $\pi$ that cannot be strictly improved upon in terms of both expected improvement and $z$-score.
2.2 Convex Relaxation Framework
This frontier is difficult to characterize directly as one of the objectives (the expected $z$-score) is not convex. We start by showing that a superset of $\Pi_{\mathrm{Opt}}$ is given by a family of convex optimization problems $\Pi_{\Lambda}$. Then, we decompose the first order condition for this family of problems in order to characterize $\Pi_{\Lambda}$, which allows us to characterize $\Pi_{\mathrm{Opt}}$. Proofs for all claims are given in Appendix 8. {repeatproposition} We have $\Pi_{\mathrm{Opt}}\subseteq\Pi_{\Lambda}$, where $\tau_{\text{max}}=\max_{\pi}\tau(\pi)$ and
$\Pi_{\Lambda}=\bigcup_{\lambda\in[0,\tau_{\text{max}}]}\operatorname*{\arg\min}_{\{\pi\in(\Delta^{K})^{N}|\tau(\pi)=\lambda\}}s^{2}(\pi).$